A bond's credit rating provides a guide to its price. Assume Aaa bonds yield 4.8% and Baa bonds yield 5.8%. Assume a 10% five-year bond with annual coupons and a face value of $1,000. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

a. What is the bond's price if it is rated as Aaa?

Bond price $

b. What is the bond's price if it is rated as Baa?

Bond price $
Answer #1

a. What is the bond's price if it is rated as Aaa?

BEGYKP8HycMnpCUFruAAAAAASUVORK5CYII=

PV factor working

=(1-(1+r)^n-1)/r

=+(1-(1+4.8%)^(-5))/4.8%=4.3535

(1+r)^n/r

=+1/(1+4.8%)^5=.7910

b

b. What is the bond's price if it is rated as Baa?

wNz2NPMgEZYcgAAAABJRU5ErkJggg==

PV factor working

=(1-(1+r)^n-1)/r

=+(1-(1+5.8%)^(-5))/5.8%=4.2354

(1+r)^n/r

=+1/(1+5.8%)^5=.7543